OpenStudy (anonymous):

Solve this system of equations by graphing. First graph the equations, and then type the solution. x–3y=-6 x+3y=-6

OpenStudy (hlambach):

It will be much easier if you first isolate y. Can you do that?

OpenStudy (anonymous):

yea so x-y=-2 & x+y=-2

OpenStudy (anonymous):

@hlambach

OpenStudy (hlambach):

No, as in find what y is. So the first one would be: x–3y=-6 -x -x -3y= -x - 6 /-3 /-3 y = 1/3x + 2 Can you do the second one?

OpenStudy (hlambach):

Do you see a ton of question marks, or is that just me? If yes, then the only thing it is supposed to be is the original equation. If not, don't mind this.

OpenStudy (anonymous):

i do

OpenStudy (hlambach):

x-3y = -6 -x -x -3y= -x - 6 /-3 /-3 y = 1/3x + 2

OpenStudy (hlambach):

But can you do the second one?

OpenStudy (anonymous):

x+3y=-6 -x -x 3y=-6-x /-3 /-3 y=2-x or y=2-1/1?

OpenStudy (anonymous):

is that right?

OpenStudy (hlambach):

Remember that when there is only an x, it actually means 1x, so then you divide the 1 by 3, which would then be y=-1/3x + 2

OpenStudy (hlambach):

So you have your two formulas y = 1/3x + 2 y = -1/3 + 2 Do you know how to graph that?

OpenStudy (hlambach):

@Sara_adalla